OpenStudy (anonymous):

simplify cos60deg cos30deg-sin60degsin30deg

OpenStudy (anonymous):

0

OpenStudy (lgbasallote):

hint \[\cos ( a + b) = \cos a \cos b - \sin a \sin b\] does that help?

OpenStudy (anonymous):

yes but i got another one involving vector if you got a moment

OpenStudy (lgbasallote):

that seems very distant from this topic...

OpenStudy (anonymous):

never mind the vector could you give me some more information on simplifying the above expression

OpenStudy (anonymous):

if u dont kne that formula ...just subs values of cos 60 and cos 30 and sin

OpenStudy (lgbasallote):

like i told you...use that formula i gave...for example \[\cos(20) \cos (40) - \sin(20)\sin (40) \implies \cos (20 + 40) \implies \cos (60)\]

OpenStudy (anonymous):

ok i would like to say thank you and move on to a harder question, sin2A/1+cos2A

OpenStudy (callisto):

I would like to say please start a new post for a new question, thanks!

OpenStudy (anonymous):

ok